OpenStudy (anonymous):

The length of a rectangle is 4cm more than breadth,the area of a rectangle is 96cm^2.Find the sides (ANSWERS: 8cm and 12cm)

OpenStudy (anonymous):

where r u stuck?

OpenStudy (anonymous):

i did'nt understand the Q

OpenStudy (unklerhaukus):

\[A_{\text{rectangle}}=l\times b\] \[l=b+4\]

OpenStudy (unklerhaukus):

\[A_{\text{rectangle}}=(b+4)\times b\] \[96=b^2+4b\] \[b^2+4b-96=0\] \[b=\frac{-(4)\pm\sqrt{(4)^2-4\times(1)\times(-96)}}{2(1)}\]

OpenStudy (unklerhaukus):

you will get solutions for b , only one of them can be the breadth of a rectangle, use this to find the length,